OpenStudy (anonymous):

What is the molar mass of a gas if 16.0 L of the gas has a mass of 19.0 g at 765 torr and 20.0 °C?

OpenStudy (mos1635):

P*V=n*R*T n=m/Mr ...............................

OpenStudy (anonymous):

ok..i will work it out and see what i get..thank you

OpenStudy (anonymous):

what is Mr?

OpenStudy (mos1635):

what you are using for R??

OpenStudy (mos1635):

note: convert units to mach units of R

OpenStudy (anonymous):

r is 0.0821 no?

OpenStudy (chmvijay):

8.314 j / mol /K

OpenStudy (anonymous):

so i did (765x16)=n(0.0821x293.15)

OpenStudy (chmvijay):

find n now

OpenStudy (anonymous):

n24.07 so i divided to cancel 24.07

OpenStudy (anonymous):

divided( 765x16)/24.07

OpenStudy (anonymous):

508.51

OpenStudy (anonymous):

taht's my moles soo...i did 508.51/19.0

OpenStudy (chmvijay):

R= 8.314 u have not took from where u took R value

OpenStudy (mos1635):

convert torr to N/m3 convert lt to m3

OpenStudy (mos1635):

Mr=molar mass

OpenStudy (mos1635):

correction:convert torr to N/m2

OpenStudy (chmvijay):

@mos1635 help the poor @Cardonuts614 as he is having final tomorrow bye good night

OpenStudy (anonymous):

i get 26.7...but the right anwer is 28.4

OpenStudy (anonymous):

i'm a she :-P

OpenStudy (mos1635):

if R=0,082...lt*atm/mole*K then convert torr to atm and lt to lt (no need)

OpenStudy (mos1635):

if R=8,314..joule/mole*K convert torr to N/m2 convert lt to m3

OpenStudy (mos1635):

what R you are using?

OpenStudy (anonymous):

ok that's where i 'm confused...why is R = 8.314?? teacher told me it was 0.0821Latm/Kmol

OpenStudy (mos1635):

deferant units no matter

OpenStudy (mos1635):

ok let's go with R=0.082 L*atm/K*mol

OpenStudy (mos1635):

convert torr to atm

OpenStudy (anonymous):

ok..give me a min and i'll try it again

OpenStudy (mos1635):

what are you getting at?

OpenStudy (anonymous):

i'm stuck...n=pv/rt right?

OpenStudy (chmvijay):

LOL yes

OpenStudy (mos1635):

ok

OpenStudy (mos1635):

\[\frac{ m }{ M }=\frac{ P*V }{ R*T }\]

OpenStudy (anonymous):

.6714

OpenStudy (anonymous):

that's what i got for n

OpenStudy (mos1635):

n=m/Mr Mr=m/n ......

OpenStudy (anonymous):

got it!

OpenStudy (mos1635):

19/0.6714=28.299...

OpenStudy (anonymous):

the answer on the quiz was 28.4 but it's close enough

OpenStudy (mos1635):

yes some times the deferences are due to how many decimals we are using....

OpenStudy (anonymous):

that's right..

OpenStudy (mos1635):

good luck with your final BB

OpenStudy (anonymous):

i'll need the luck. thank you.